Find indicated quantity if it exists.

La x + 3 if x < -2 54. Let f(x) Find IVx+ 2 if x > -2 (A) lim+ f(x) (B) lim-f(x) (C) lim, f(x) (D) f(-2) x-27
Answer #1

fexna S Xt3 [x+ 2 { af x 4.2 is X7-2 (A) 0 5-2 +2 Le lim fixo = X-2t lim Sx42 X4-24 (B) lim X+3) (-2+3) 1 L = lim fix) x720 X